Question #17: You just started your Retirement Savings Plan by investing $5,000. You do not plan on making any additional investments in the Plan.

Required: If you can earn 8.4% on this investment, what will your Retirement Savings Plan be worth in

a) 35 years?

b) 40 years?
Answer #1

Future value=Present Value*(1+r)^t

1.
=5000*1.084^35=84138.5225046972

2.
=5000*1.084^40=125933.505450691